### v3

# The impact of kinetic and global effects on ballooning 2nd stable pedestals of conventional and low aspect ratio tokamaks

Abstract
The EPED model [P.B. Snyder et al 2011 Nucl. Fusion 51 103016] had success describing the pedestals of the Type-I ELM and QH-mode operations in conventional tokamaks, by combining kinetic ballooning mode (KBM) and peeling-ballooning (PB) constraints. Within EPED, the KBM constraint is usually approximated by the ideal ballooning mode (IBM) stability threshold. It has been noted that quantitative differences between local ideal MHD and gyro-kinetic (GK) ballooning stability can be larger at low aspect ratio. KBM critical pedestals are consistent with observations in initial studies on conventional and spherical tokamaks. In this work, the application of a reduced model for the calculation of the kinetic ballooning stability boundary is presented based on a novel and newly developed Gyro-Fluid System (GFS) code [G.M. Staebler et al 2023 Phys. Plasmas 30 102501]. GFS is observed to capture KBMs in DIII-D as well as the NSTX(-U) pedestals, opening a route integrating this model into EPED. Finally, high-$n$ global ballooning modes are observed to limit the access to the local 2nd stability and thus provide a transport mechanism that constrains the width evolution with $\beta_{p,ped}$. The high-$n$ global ballooning stability is approximated by its ideal MHD analogue using ELITE. It is shown that nearly local high-$n$ with $k_{y}\rho_{s}\sim 1/2$ modes can provide a proxy for the critical $\beta_{p,ped}$ when a 2nd stable access exists on DIII-D plasmas. The use of GFS and ELITE scaling in EPED provided an improved agreement in comparison to EPED1 with DIII-D pedestal data.
1 Introduction
As the efforts for the design and construction of a tokamak Fusion Pilot Plant (FPP) becomes more intense, the need for reliable predictive modeling capabilities becomes imperative as current tokamak devices operate far from a reactor parameter space. This requires deep physics understanding for the creation of robust reduced models that are accurate but also fast enough to be integrated in reactor design and optimization tools.
The general goal of such design loops is the creation of FPPs that have good enough core performance in order for fusion reactions to take place. The fusion performance of the tokamak is often measured in terms of the triple product, $nT\tau_{E}$, where $n$ and $T$ is the density and temperature of the deuterium and tritium mixture and $\tau_{E}$ is the energy confinement time. In current tokamak devices, the maximization of the core pressure is often achieved by the formation of an edge transport barrier that leads to the formation of a pedestal on which the core pressure sits. Although this operational regime, called the H-mode, is beneficial for performance it is accompanied by the destabilization of edge localized modes (ELMs). In a reactor, just a few large ELMs will lead to excessive sputtering, erosion and melting of the divertor components and therefore it is necessary to understand, predict and eventually avoid the occurrence of such events.
For the design of a tokamak FPP, the pedestal conditions are extremely important. The core of the plasma will be most likely close to turbulence critical gradient limits to achieve significant fusion gain and therefore the height of the pedestal is one of the main drivers to maximize performance. In addition, continuous FPP operation highly relies on non-inductive current and the pedestal will contribute significantly to the bootstrap current fraction. As a result, accurate pedestal modeling capabilities are necessary for predicting performance and stability limits in order to develop strategies for small/no ELM and non-inductive operation.
The most successful model to date used to predict the pedestal structure and the onset of large type-I ELMs and QH-mode is the EPED model [1]. EPED is based on two simple ideas that define how the pedestal, which is characterized by the width $\Delta_{\psi_{N}}$ and height $\beta_{p,ped}$, evolves. That relation is defined as $\Delta_{\psi_{N}}=c_{1}\beta_{p,ped}^{c_{2}}$. Here $\beta_{p,ped}=2\mu_{0}p_{ped}/(\mu_{0}I_{p}/L)^{2}$, where $p_{ped}$ is the top of the pedestal pressure, $I_{p}$ is the plasma current and $L$ is the circumference of the plasma boundary. There exists a local “stiff” transport constraint that limits the pressure gradient of the pedestal and a global constraint based on ideal MHD instabilities that limits the maximum width and height of the pedestal. The pedestal width and height (pressure gradient) evolves close to the kinetic ballooning limit, and type-I ELMs or QH-mode occur when eventually the peeling-ballooning modes become unstable. Implicitly the relation between the width and height has profound implications for the destabilization of peeling-ballooning modes. Therefore it is important to be able to capture this relation as accurate as possible.
In conventional aspect ratio tokamaks it is experimentally observed that $c_{1}\sim 0.06-0.13$ and $c_{2}\sim 0.5$ [2][3][4][5][6][7] , which is also theoretically reproduced considering the stability of local ideal or kinetic ballooning modes. The same square root dependance of the width with $\beta_{p,ped}$ is also observed in MAST with somewhat larger $c_{1}=0.12-0.15$ [8]. However, in NSTX a different relation is observed experimentally with $c_{1}\sim 0.4$ and $c_{2}\sim 1$ [9]. Considering the ideal ballooning limit of NSTX pedestals, almost a linear relation $c_{2}\sim 0.8-1$ between the width and the height is reproduced by the ideal ballooning mode (IBM) limit but the $c_{1}\sim 0.2-0.3$, which implies a much larger $\beta_{p,ped}$ for a certain $\Delta_{\psi_{N}}$ than what is measured. This discrepancy has been recently attributed to the difference between kinetic and ideal ballooning mode threshold [10] which seems to be enhanced in low aspect ratio tokamaks. The impact of kinetic effects on the critical $\beta_{p,ped}$ results in a more strict stability limit (lower critical $\beta_{p,ped}$) and therefore a larger $c_{2}$ coefficient.
Moreover, one of the features of local ideal and kinetic ballooning stability is the existence of a 2nd stable region at high enough $\beta^{\prime}=2\mu_{0}/B^{2}dp/dr_{N}$, where $r_{N}=r/a$ and $a$ is the plasma minor radius. In cases of low magnetic shear $\hat{s}=d\ln{q}/d\ln{r}$ and shaped magnetic geometry no stability boundary exists with increasing $\beta^{\prime}$. This feature makes the determination of the width and height constraint impossible based on local ballooning physics alone. It remains an open question whether local analysis is appropriate in the pedestal region where the width of the pedestal is often smaller than the $k_{y}\rho_{s}$ of the most unstable kinetic ballooning mode (KBM). It has been shown through global gyro-kinetic simulations that kinetic effects in fact close the access to a 2nd stable region at low magnetic shear [11]. However, the gyro-kinetic ordering is intrinsically local, excluding global equilibrium variations. The impact of finite-$n$ effects on 2nd stability access was also examined considering ideal MHD physics based on the interplay between local high-$n$ kink/peeling and ballooning modes [12][13] introducing global equilibrium variations.
In this work, we attempt to estimate $\Delta_{\psi_{N}}=c_{1}\beta_{p,ped}^{c_{2}}$ using a novel and newly developed Gyro-Fluid System (GFS) code [14] that has successfully captured KBMs on NSTX. GFS defines a gyro-kinetic fluid moment system of equations that capture the important kinetic effects but at far less computational cost than existing gyro-kinetic codes, making the integration with frameworks like EPED possible. In addition, using the ideal MHD code ELITE [15] the global effects on 2nd stability are examined and observed to provide a realistic width scaling for DIII-D plasmas. Although ELITE does not contain kinetic physics, the computational expense of global gyro-kinetic simulations prohibits their use for such iterative predictive studies. Combining the pedestal scaling as obtained from GFS and ELITE with EPED better agreement is observed in comparison to EPED1 on a set of DIII-D plasmas.
This paper is organized as follows. Section.2 introduces the GFS model and the procedure that is used to identify the critical $\beta_{p,ped}$ and obtain the width scaling $\Delta_{\psi_{N}}=c_{1}\beta_{p,ped}^{c_{2}}$. Section.3 presents the predicted width scaling on DIII-D and NSTX plasmas and examines the impact of geometry and plasma parameters on NSTX #139047 local pedestal stability. Section.4 presents the impact of global effects on local ideal ballooning 2nd stable DIII-D pedestals and the use of high-$n$ peeling-ballooning modes as a constraint for the evolution of the width and height. Section.5 demonstrates the applicability of GFS and ELITE as tools for the calculation of the KBM constrain using EPED predictions on DIII-D pedestals. Section.6 discusses and summarizes the results.
DIII-D NSTX
pulse # 131999 139047
$R_{0}$ [m] 1.695 0.863
$r$ [m] 0.579 0.599
$\kappa$ 1.807 2.312
$\delta$ 0.148 0.536
$B_{T}$ [T] 2.098 0.486
$I_{p}$ [MA] 1.487 0.965
$\beta_{N}$ 1.851 4.778
[…]
$n_{e,ped}/n_{e,sep}$ 4.0 1.6
$n_{e,0}/n_{e,ped}$ 1.5 1.5
$T_{e,sep}$ [eV] 75 80
Table 1: Table with global plasma parameters for DIII-D #131999 and NSTX #139047 plasmas that are used to create bootstrap consistent equilibria and identify the KBM threshold using GFS.
2 Creation of ballooning critical pedestals using GFS
As it is highlighted in Section.1, kinetic effects can be important for the determination of the critical height $\beta_{p,ped}$ for a fixed width $\Delta_{\psi_{N}}$. It is well understood that ion drift resonances and trapped particles can further destabilize the kinetic analogue of the ideal ballooning mode close to marginal stability [16][17][18]. Those effects are enhanced in low aspect ratio and strongly shaped plasmas, which creates the need for an accurate but fast model to capture those effects. The recent development of the GFS code opens the road for a quick and accurate model to be coupled with integrated modeling tools.
GFS solves the fluid (velocity) moments of the linear gyro-kinetic equation using an orthogonal set of basis functions that reduce the computational complexity of the system. The full electromagnetic system considering perturbations on the electrostatic potential $\delta\phi$, parallel vector potential $\delta A_{\parallel}$ and parallel magnetic field $\delta B_{\parallel}$ is included through quasi-neutrality and Ampere’s law. Pitch angle scattering and like-species collisions are retained in the linearized collision operator. The basis functions used are Hermite polynomials in the parallel to the field-line and parallel velocity coordinates and Laguerre polynomials for the perpendicular velocity coordinate. A more detailed description of the GFS equations is given in [14].
a) Refer to caption
b) Refer to caption
Figure 1: a) The $\hat{s}-\alpha$ diagram for the IBM stability boundary and KBM stability boundary as calculated from GFS. b) Comparison of normalized growth rate $\gamma/\omega_{s}$ between GFS and CGYRO with varying temperature gradient normalized length scale $L_{T}$ for $k_{y}\rho_{s}=0.1$.
In order to test the applicability of the GFS code in capturing KBMs, a large set ($\sim 800$ cases) of NSTX KBMs using CGYRO [19] have been compared. More details on this benchmark exercise can be found in [20][21], where GFS reproduced the linear frequency and growth rates of those KBM cases well even at low velocity space resolution. Furthermore, GFS is used to obtain the local $\hat{s}-\alpha$ KBM stability boundary and compared to the IBM stability boundary. The normalized pressure gradient is defined as $\alpha=-\mu_{0}/(2\pi^{2})dV/d\psi\sqrt{V/(2\pi^{2}R_{0})}dp/d\psi$, where $V$ is the volume and $R_{0}$ the major radius. . As it can be observed from Figure(1) in the limit where $k_{y}\rho_{s}\rightarrow 0$ (or $\omega_{p}\rightarrow 0$) the KBM reduces to the IBM, where $k_{y}$ is the binormal wavenumber, $\rho_{s}$ is the main species Larmor radius and $\omega_{p}$ the diamagnetic drift frequency. When finite $k_{y}\rho_{s}$ effects are introduced it can be observed that a stricter limit is observed as the KBM gets further destabilized by the drifts. Finally, the stability boundary can be found by computing at which value of $\beta_{p,ped}$ the transition from an electrostatic to an electromagnetic mode occurs. Due to the “stiffness” of KBMs with pressure gradient, at low $k_{y}\rho_{s}\sim 0.1-0.05$ a sharp increase in the growth rate signifies the onset of such modes. Such an example is given in Figure(1) where a comparison between GFS and CGYRO is given. Even at low parallel $n_{u}$ and perpendicular $n_{e}$ velocity resolution a good agreement is observed between the two codes for the normalized growth rate $\gamma/\omega_{s}$, where $\omega_{s}=c_{s}/a$ with $c_{s}$ being the sound speed. The use of low $k_{y}\rho_{s}$ minimizes ion/electron temperature gradient (ITG/ETG) and trapped electron (TEM) instabilities while the low velocity resolution is observed to capture criticality at a satisfactory level while stabilizes micro-tearing modes (MTMs). In such a way the KBM onset is effectively computed.
[…]
3 Predictive pedestal width and height scaling using GFS
a) Refer to caption
b) Refer to caption
Figure 2: a) Comparison of the pedestal width scaling with $\beta_{p,ped}$ between DIII-D #131999 and NSTX #139047 and b) $\hat{s}-\alpha$ kinetic ballooning stability boundary at the middle of the pedestal between DIII-D and NSTX plasmas. The solid lines represent the ideal ballooning stability boundary.
a) Refer to caption
b) Refer to caption
Figure 3: a) Comparison of the pedestal width $\Delta_{\psi_{N}}$ scaling with $\beta_{p,ped}$ and b) the $\hat{s}-\alpha$ stability boundary at the middle of the pedestal considering different toroidal magnetic field $B_{T}$ on the geometric axis for NSTX #139047.
a) Refer to caption
b) Refer to caption
Figure 4: a) Comparison of the pedestal width $\Delta_{\psi_{N}}$ scaling with $\beta_{p,ped}$ and b) the $\hat{s}-\alpha$ stability boundary at the middle of the pedestal considering different aspect ratio AA for NSTX #139047.
a) Refer to caption
b) Refer to caption
Figure 5: a) Comparison of the pedestal width $\Delta_{\psi_{N}}$ scaling with $\beta_{p,ped}$ and b) the $\hat{s}-\alpha$ stability boundary at the middle of the pedestal considering different elongation $\kappa$ for NSTX #139047.
a) Refer to caption
b) Refer to caption
Figure 6: a) Comparison of the pedestal width $\Delta_{\psi_{N}}$ scaling with $\beta_{p,ped}$ and b) the $\hat{s}-\alpha$ stability boundary at the middle of the pedestal considering different triangularity $\delta$ for NSTX #139047.
a) Refer to caption
b) Refer to caption
Figure 7: Pedestal width $\Delta_{\psi_{N}}$ and height $\beta_{p,ped}$ scaling over a wide range of plasma discharges from a) DIII-D and b) NSTX using the GFS code. The colored lines indicated the particular pulse while the black line indicates the average trend.
For each equilibrium every location in the pedestal defines a “width” and has an associated critical gradient $d\beta_{p,ped}/d\Delta_{\psi_{N}}$. Considering a smaller number of equilibria with different $\Delta_{\psi_{N}}$, a large number of local widths and heights are obtained such that a robust scaling is calculated. For each equilibrium GFS is used to perform a local scan of increasing $d\ln T/dr_{N}$ or $\beta$ (i.e. increasing $\beta^{\prime}$) to find the critical value. The equilibrium $T_{e,ped}$ is then iterated until the critical and equilibrium $\beta_{p,ped}$ are the same within a small relative error.
The application of GFS as a tool to predict the pedestal width is examined by considering ELMy H-mode plasmas for DIII-D and NSTX with the ultimate goal being the determination of $\Delta_{\psi_{N}}=c_{1}\beta_{p,ped}^{c_{2}}$. Table(1) shows the plasma parameters for the different tokamak configurations. As it can be seen in Figure(2), the predicted $\Delta_{\psi_{N}}=c_{1}\beta_{p,ped}^{c_{2}}$ for those three cases is in agreement with the existed empirical scalings. In DIII-D #131999 $c_{1}=0.09$ and $c_{2}=0.48$. The NSTX #139047 scaling follows Diallo et al [9] with $c_{2}=0.38$ and $c_{2}=0.98$ demonstrating the applicability of GFS as a tool to calculate the KBM critical $\beta_{p,ped}$ in both medium and low aspect ratio tokamaks and wide range of plasma parameters.
There exists a striking deviation in the scaling between NSTX(-U) and DIII-D and we can identify several differences between those cases. Although DIII-D is a medium aspect ratio tokamak, both NSTX and MAST are low aspect ratio tokamaks and have different scaling, so the difference in scaling cannot be solely due to the aspect ratio. In this cases, the deviations between DIII-D #131999 and NSTX #139047 are found in the toroidal field $B_{T}$, plasma current $I_{p}$ and shaping (elongation $\kappa$, triangularity $\delta$) leading to differences in the $q$-profile and qualitative differences on the characteristics of the pedestal with respect to the $\hat{s}-\alpha$ diagram and access to 2nd stability as it can be seen in Figure(2).
In order to understand the main drivers for the difference in scaling between these cases on DIII-D and NSTX(-U), the equilibrium parameters of the NSTX #139047 plasma are varied individually to identify their impact. Initially, the toroidal magnetic field $B_{T}$ is varied by $\pm 15\%$. It is interesting to note that higher $B_{T}$ increases the $q$-profile, elevates the $\hat{s}-\alpha$ stability boundary and leads to easier access to the 2nd stable region, as can be seen from Figure(3). At higher $B_{T}$ the pedestal approaches the 2nd stable region and a strong reduction of the exponent $c_{2}$ is observed. Geometrical effects are observed to play a significant role on pedestal performance. The low aspect ratio $A$ plays an important role on enhancing pedestal performance, even though the pedestal is still limited by the 1st stability boundary as can be seen in Figure(4). Higher values of elongation $\kappa$ leads to larger pedestal height $\beta_{p,ped}$ for a fixed width $\Delta_{\psi_{N}}$ as can be seen from Figure(5). Higher triangularity $\delta$ is also observed to have a beneficial impact on the pedestal height $\beta_{p,ped}$ as can be seen from Figure(6). Although, due to the large magnetic shear $\hat{s}$ the middle of the pedestal is still far from the “nose” of the stability boundary even with strong shaping. The parametric dependence of NSTX #139047 on geometry is in good agreement with Ref.[24]. Using GFS the width scales as $\kappa^{-1.86}$, $0.67^{\delta}$ and $A^{1.67}$, while in Ref.[24] $\kappa^{-1.8}$, $0.5^{\delta}$ and $A^{1.5}$.
Access to 2nd stability, which is typically the case for DIII-D pedestals, is found to correlate with a lower $c_{2}$ exponent. The arguments made by Ref.[2] are followed to understand the impact of accessing 2nd stability. The pedestal width can be approximated by $\Delta_{\psi_{N}}\sim\beta_{p,ped}/\alpha_{crit}$. Close to the “nose” of the stability diagram $\alpha_{crit}\sim\hat{s}^{-1/2}$ is satisfied and due to the bootstrap dominated current $\hat{s}\sim 1/<j>\sim 1/\beta_{p,ped}$. As a result, substituting those relations to the width we find $\Delta_{\psi_{N}}\propto\beta_{p,ped}^{1/2}$. However, at high $\hat{s}$ where $\alpha_{crit}$ is constant the width is simply $\Delta_{\psi_{N}}\propto\beta_{p,ped}$. It can be concluded that the proximity and access to 2nd stability are crucial for understanding the differences in pedestal width/height scaling. It is interesting to note that at low shaping access to 2nd stability might be closed due to the strong coupling of edge localized kink/peeling and ballooning modes leading to decreasing $\alpha_{crit}$ as $\hat{s}$ reduces. However, such equilibrium evolution is unphysical due to the connection between $\hat{s}$ and bootstrap current, providing a natural barrier for the pedestal evolution.
Finally, considering a number of cases from DIII-D and NSTX, GFS is used to compute the pedestal width/height scaling. As it can be observed from Figure(7) an excellent agreement is observed statistically for both DIII-D, with $\Delta_{\psi_{N}}=0.08\beta_{p,ped}^{0.48}$, and NSTX, with $\Delta_{\psi_{N}}=0.43\beta_{p,ped}^{1.01}$. It is interesting to note in all those cases, DIII-D pedestals are found close to the nose of the $\hat{s}-\alpha$ diagram, while NSTX is ballooning limited at high $\hat{s}$ where $\alpha_{crit}$ is roughly constant. This qualitative difference in $\hat{s}-\alpha$ stability is identified as the main difference in the observed scaling.
4 Global effects on local 2nd stable pedestals
Figure 8: Comparison of the $c_{1}$ coefficient for $\Delta_{\psi_{N}}\propto\beta^{1/2}_{p,ped}$ scaling between GFS, ELITE and BALOO using a set of DIII-D plasma discharges.
a) Refer to caption
b) Refer to caption
Figure 9: Comparison of a) the pedestal pressure ($P=2n_{e}T_{e}$) and b) the width $\Delta_{\psi_{N}}$ between EPED1 (EXP/ELITE), EPED1.63 (BALOO/ELITE), EPED2 (ELITE/ELITE) and EPED3 (ELITE/GFS) with DIII-D experimental data.
It becomes apparent from Section(3) that DIII-D pedestals can access the 2nd stable region under certain conditions. In fact in various experimental cases, it is observed that plasma parameters do exist where the pedestal is ideal ballooning 2nd stable. Therefore, if both local IBMs and KBMs are found stable another mechanism should exist that limits the pressure gradient in those cases. ITGs, ETGs and MTMs, if unstable, do not necessarily limit the pressure as they are driving predominantly thermal transport and the pedestal pressure gradient can keep evolving through density built up. TEMs can certainly be found unstable in the pedestal but not enough information exists on their impact as a pedestal constraint. Global gyro-kinetic simulations in JET demonstrated that global effects lead to unstable KBMs in 2nd stability where local KBMs are stabilized [11]. However, performing global gyro-kinetic simulations to examine this effect is impractical due to the immense computational expense. An alternative approach relies on the global ideal MHD model, that is found to provide a destabilizing mechanism through the coupling of kink/peeling modes and finite-$n$ effects on ballooning modes in 2nd stability [12][13].
To demonstrate this, a set of DIII-D plasma discharges are considered as test cases and the ELITE code is used to calculate the stability of high-$n$ modes. When $n\sim(\rho^{*}q)^{-1}$, where $\rho^{*}=\rho_{s}/r$, the global mode is nearly local in nature as it is localized solely in the steep gradient of the pedestal. In this analysis, modes with $k_{y}\rho_{s}<1/2$ are assumed to be non-local as their extent can be larger than the pedestal, while modes with $k_{y}\rho_{s}>1/2$ will be stabilized by finite ion Larmor radius effects and are therefore not considered. ELITE is used to find the critical $\beta_{p,ped}$ where $k_{y}\rho_{s}=1/2$ modes are marginally unstable. It is interesting to note that the kink/peeling contribution is necessary for the instability even at such high-$n$, as the ballooning drive alone is not enough to overcome the stabilizing field-line bending. For typical DIII-D pedestals, it is found that $n\sim 40-60$ depending on the pedestal temperature and $q$-profile ($\rho^{*}$, $r$ and $q$ are taken at the middle of the pedestal) satisfy $k_{y}\rho_{s}=1/2$.
The DIII-D cases under consideration are local KBM limited based on GFS calculations, but have partial or complete access to ideal ballooning 2nd stability using BALOO [25]. Those case are used as a demonstration to compare the $c_{1}$ coefficient as computed for ELITE limited pedestal using $k_{y}\rho_{s}=1/2$ with GFS and BALOO limited pedestals. The exponent of $\beta_{p,ped}$ is fixed to $c_{2}=1/2$ in this exercise. As it can be seen by Figure(8) GFS and ELITE result is in good agreement, while the scaling using BALOO deviates from the kinetic and finite-$n$ result. This has profound implications for EPED calculations as the width/height scaling has a strong impact on the maximum achievable pedestal pressure.
5 Application of kinetic and finite-$n$ effects on EPED
To further demonstrate the application of this approach, a number of DIII-D pedestal cases are examined with EPED using GFS, ELITE and BALOO as the KBM constraint. It should be mentioned that in the majority of the pedestals under consideration, ideal stability suggests that significant part of the pedestal has 2nd stability access which is limited by kinetic or finite-$n$ effects. As a result,t the BALOO calculation might not provide an accurate scaling. For each plasma pulse GFS, ELITE and BALOO are used to evaluate the pedestal width/height scaling, which is then used in EPED to compute the global MHD boundary with ELITE. The standard scaling of EPED1, where $\Delta_{\psi_{N}}=0.076\beta_{p,ped}^{1/2}$, is also compared. As it can be observed from Figure(9) better agreement is observed for both pedestal pressure $P_{ped}=2n_{e,ped}T_{e,ped}$ and width $\Delta_{\psi_{N}}$ between EPED1 (EXP/ELITE), EPED1.63 (BALOO/ELITE), EPED2 (ELITE/ELITE) and EPED3 (GFS/ELITE) with experimental measurements when using the global or kinetic constraint width/height. In particular, the standard deviation of EPED1 is $\sigma_{P}=0.14\%$ and $\sigma_{\Delta}=0.21\%$, for EPED1.63 is $\sigma_{P}=0.17\%$ and $\sigma_{\Delta}=0.26\%$, for EPED2 is $\sigma_{P}=0.12\%$ and $\sigma_{\Delta}=0.2\%$, while for EPED3 is $\sigma_{P}=0.08\%$ and $\sigma_{\Delta}=0.15\%$. It can be concluded that using GFS or ELITE to compute the KBM constraint led to better agreement with experimental measurements for both pedestal pressure and width, highlighting the importance of kinetic or finite-$n$ effects close to 2nd stability.
6 Summary
This work focused on calculating the KBM pedestal width $\Delta_{\psi_{N}}$ and height $\beta_{p,ped}$ considering that typically takes the form of $\Delta_{\psi_{N}}=c_{1}\beta_{p,ped}^{c_{2}}$. The required improvements are motivated by the unique characteristics of NSTX(-U) pedestals, that are typically not described by the IBM stability, and the recent resolution of the discrepancy attributed to kinetic effects. The development of the GFS code allowed for an accurate and fast numerical algorithm to perform KBM stability calculations as it is shown through this work.
Experimentally relevant pedestal width/height scaling is reproduced considering the KBM critical $\beta_{p,ped}$ as calculated from GFS for a number of DIII-D and NSTX plasmas. A comparison between DIII-D and NSTX showed that the DIII-D pedestals are close to the nose of the $\hat{s}-\alpha$ stability diagram, while NSTX pedestals are found at higher $\hat{s}$ in the 1st stability boundary for the cases under consideration. Through changes on the plasma parameters and geometry of NSTX #139047, the originally calculated linear width/height scaling approached the square root dependence when the pedestal approached the nose of the stability diagram.
When the pedestal gets access to the local IBM or KBM 2nd stability region, a transport mechanism should exist that constrains the evolution of the pressure gradient. Global equilibrium effects have been identified to provide a mechanism for the destabilization of ballooning modes for local 2nd stable pedestals. In order to capture such global effects, ELITE is used as a KBM proxy for capturing unstable $k_{y}\rho_{s}\sim 1/2$ modes localized within the pedestal region to compute a critical $\beta_{p,ped}$. The coupling of kink and ballooning drive is essential and results in unstable modes even at high-$n$ when local ballooning modes are 2nd stable.
Finally, considering a set of DIII-D pedestals where partial or complete 2nd stable access is observed, the predicted electron pedestal pressure and width as computed from EPED using the GFS or ELITE scaling are in good agreement with the experimental measurements, providing a systematic approach for computing the pedestal scaling when local ballooning modes are ideal 2nd stable.
Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, Theory Program and the NSTX Research Program, using the NSTX Fusion Facility, a DOE Office of Science user facility. We thank the NSTX experimental team for providing the data analyses.
DISCLAIMER: This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise, does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
Notice: This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-public-access-plan).
The data that support the findings of this study are available from the corresponding author upon reasonable request.